A detailed solution i.e with working for the below question:

Q: For the production function Q=20K^0.5 L^0.5, determine four combination of capital and labor that will produce 100 and 200 units of output. Plot these points on a graph and use them to sketch the 100 and 200 unit isoquants.

Ans: If the production function is Q=20K^0.5 L^0.5, then four combination of capital and labor that will produce 100 and 200 units of output will be (K;L):
Q = 100 units: (1;25), (25;1), (6.25;2), (2;6.25);
Q = 200 units: (1;100), (100;1), (25;2), (2;25).
1
Expert's answer
2015-07-09T00:00:42-0400
If the production function is Q=20K^0.5 L^0.5, then four combination of capital and labor that will produce 100 and 200 units of output will be (K;L):
K = Q^2/400L, so if Q = 100, then K = 25/L
Q = 100 units: (1;25), (25;1), (6.25;2), (2;6.25);
Because, for example, 20*1^0.5*25^0.5 = 20*1*5 = 100 units.
K = Q^2/400L, so if Q = 200, then K = 100/L
Q = 200 units: (1;100), (100;1), (25;2), (2;25).